OpenStudy (anonymous):

Find x intercept y = x^2 + 2x + 1

myininaya (myininaya):

set y=0

myininaya (myininaya):

(x+1)^2=0 x=-1

OpenStudy (anonymous):

y=(x+1)^2 implies x=-1, -1, so x intercept is -1

OpenStudy (anonymous):

I may be wrong, but I have the vertex being (-1,0) wouldnt the x- intercept be different than the vertex of x?

myininaya (myininaya):

sometimes

myininaya (myininaya):

this case it is

OpenStudy (anonymous):

ok thanks guys!

OpenStudy (anonymous):

u ask for x intercept only due to that ans is x=-1